### v1

# A Survey of Event Causality Identification: Principles, Taxonomy, Challenges, and Assessment

Abstract
Event Causality Identification (ECI) has become a crucial task in Natural Language Processing (NLP), aimed at automatically extracting causalities from textual data. In this survey, we systematically address the foundational principles, technical frameworks, and challenges of ECI, offering a comprehensive taxonomy to categorize and clarify current research methodologies, as well as a quantitative assessment of existing models. We first establish a conceptual framework for ECI, outlining key definitions, problem formulations, and evaluation standards. Our taxonomy classifies ECI methods according to the two primary tasks of sentence-level (SECI) and document-level (DECI) event causality identification. For SECI, we examine feature pattern-based matching, deep semantic encoding, causal knowledge pre-training and prompt-based fine-tuning, and external knowledge enhancement methods. For DECI, we highlight approaches focused on event graph reasoning and prompt-based techniques to address the complexity of cross-sentence causal inference. Additionally, we analyze the strengths, limitations, and open challenges of each approach. We further conduct an extensive quantitative evaluation of various ECI methods on two benchmark datasets. Finally, we explore future research directions, highlighting promising pathways to overcome current limitations and broaden ECI applications.
Index Terms:
Natural Language Processing, Event, Causality, Information Retrieval, Representation Learning, Knowledge Reasoning.
publicationid: pubid: $0$–$00$ 2021 IEEE
1 Introduction
As big data continues to proliferate, the channels and methods for acquiring unstructured text are constantly expanding. One of the primary challenges now lies in how to automatically extract valuable information and knowledge from these texts. This has become a significant area of research in Natural Language Processing (NLP) and Knowledge Reasoning. Events form the core content of texts, and various research directions have emerged around events, such as Event Extraction (EE) [1], Event Relation Extraction (ERE), and Event Coreference Resolution (ECR) [2]. In recent years, Event Causality Identification (ECI) has gained increasing attention as an important and challenging task [3]. As a critical subtask of ERE, ECI aims to predict whether there is a causal relationship between given events in a text. ECI has been widely applied in tasks such as question answering systems [4], information retrieval [5], event prediction [6], knowledge graph construction [7], and reading comprehension [8].
In the ECI task, events are represented by their triggers, known as "event mentions." The task then becomes determining which event mentions in a given text have a causal relationship. Figure 1 provides an example of this. ECI is vital for text understanding and decision-making, as it uncovers the causes and effects of events, helping to analyze risks and opportunities for more informed, data-driven decisions. This capability is particularly essential in domains that require complex reasoning, such as finance [9], law [10], healthcare [11], and the military [12].
Figure 1: An example of ECI. The red boxes indicate event mentions. The blue solid arrows represent intra-sentence event causalities, while the green dashed arrows represent inter-sentence event causalities.
ECI focuses solely on extracting specific types of relations—cause and effect. However, compared to the general ERE task, identifying causalities, this task is more challenging due to several factors:
  1. 1.
    Implicitness: Causal links are often implied rather than directly stated, requiring deeper contextual and semantic understanding.
  2. 2.
    Long-distance dependencies: Causalities may span multiple sentences or paragraphs, requiring models to capture distant interactions.
  3. 3.
    Complexity: Causal chains can involve multi-step links (e.g., "earthquake→collapsed→trapped"), adding complexity to identification.
  4. 4.
    Sample imbalance: In supervised settings, negative samples (non-causal pairs) dominate, challenging models to learn effectively from fewer positive cases.
  5. 5.
    Limited labeled training data: Most datasets are small, requiring models to parse semantics well and generalize causal patterns from limited annotations.
Early ECI relied on feature pattern recognition, using cues like lexical signals [13, 14, 15], temporal features [16, 17], and co-occurrences [18, 19]. With the rise of Deep Learning (DL), more advanced methods emerged, enabling models to better capture contextual information by deeply encoding text semantics [20]. Transformer-based [21] Pre-trained Language Models (PLMs) [22] have revolutionized ECI, as they are trained on large corpora, enhancing semantic understanding and producing high-quality event and context representations, thereby improving causal identification [23, 24, 25]. Since 2023, Large Language Models (LLMs) have become widely popular. Using vast datasets and large-scale self-supervised learning, LLMs have gained stronger knowledge and contextual understanding, enabling them to handle few-shot and zero-shot tasks [26, 27]. This has advanced research in event extraction, relation extraction, and question answering. However, research by Gao et al. [26] highlights that while LLMs can perform zero-shot ECI with simple prompts, they are prone to "causal hallucination," leading to many false positives. Hence, even though LLMs have significantly advanced tasks related to textual event analysis, ECI remains highly challenging.
Research on ECI is still in its early stages, but the field is expanding rapidly. Recent years have seen a notable increase in ECI-related publications, particularly between 2022 and 2024, indicating growing interest and engagement. Though still a relatively small area, ECI research is diversifying in methodological approaches and gaining visibility in high-impact venues. This upward trend suggests that ECI will continue to attract significant attention in the near future.
Several surveys have examined event causality and ERE in natural language. Asghar [28] reviewed early rule-based and statistical methods for causal extraction but did not cover deep learning (DL) approaches. Zhao et al. [29] and Xie et al. [30] provided overviews of advances in ERE, with Zhao et al. summarizing general trends in entity causality and Xie et al. comparing DL-based supervision methods. Liu et al. [31] focused on causal and temporal relationships but with limited coverage of recent methods, while Yang et al. [32] and Ali et al. [33] reviewed explicit and implicit causality extraction, limited to work prior to 2021. Liu et al. [34] covered methods, challenges, and datasets for event relation identification but lacked detailed method classification.
In this paper, we provide a comprehensive review and analysis of the current state of research on ECI. We systematically outline the core concepts, key technologies, and methodological frameworks in this field, and conduct a quantitative evaluation and comparison of existing approaches. Furthermore, we present a thorough outlook on the future development trends of this promising research area. Our main contributions are as follows:
  • We detail several concepts of ECI, including problem formalization, datasets, evaluation protocols, and key technologies.
  • We propose the first comprehensive classification framework for ECI (Figure 2), summarizing the technical features of various methods, evaluating their strengths and limitations.
  • We perform a quantitative comparison of different ECI methods using experimental data reproduced on two common datasets.
  • We discuss future directions in ECI, highlighting key challenges and potential solutions for advancing the field.
Figure 2: Taxonomy of ECI models, distinguishing SECI and DECI subtasks. For Feature Pattern-Based Matching SECI methods, only reference indices are provided due to the age of these methods and absence of model names. Other methods are labeled as "Model Name/Author Name + Reference Index" for clarity.
The remainder of this paper is organized as follows: Section 2 presents the relevant concepts of ECI, reviews commonly used datasets, and evaluation metrics. Section 3 introduces key technologies for ECI. Section 4 provides a comprehensive overview of the classification framework for ECI, summarizes core techniques of various models, and analyzes their strengths and weaknesses. Section 5 presents a quantitative evaluation of the performance of existing classical methods on two datasets. Section 6 outlines future research directions for ECI.
2 Background
In this section, we begin by defining key concepts, including events, event causality, and ECI. We then provide a comprehensive review of the benchmark corpora commonly used for extracting event causalities in recent years. Lastly, we outline the evaluation metrics employed to assess the performance of ECI models.
2.1 Definitions and Problem Formalization
2.1.1 Event
An “event” is an objective fact defined by specific individuals, objects, and actions at a certain time and place, typically stored in unstructured formats like text, images, and videos. This study focuses on events contained within textual data.
For example, the sentence "Around 8:40 AM on February 6, shortly after a subway train left a station in Moscow, the second carriage suddenly exploded, and the train caught fire." describes two events: "the train exploded after departure" and "the train caught fire." Textual events often include elements like triggers and arguments. To facilitate analysis, triggers are used as event mentions. In this example, "exploded" and "fire" serve as the trigger words for the respective events.
2.1.2 Event Causality
Broadly speaking, causalities between events refer to the driving connections between them. If two events, A and B, have a causality represented as A→B, it means that the occurrence of A is likely to lead to the occurrence of B, with A being the cause and B being the effect. In this context, the causality between events in a text can be defined as follows:
Definition 1: Event Causality. In a given text $\mathcal{D}$, if events $e_{1},e_{2}\in\mathcal{D}$ have a causality $e_{1}\rightarrow e_{2}$, it implies that $e_{1}$ directly or indirectly leads to the occurrence of $e_{2}$, while $e_{2}$ does not lead to $e_{1}$.
Example 1:
Text: "A car bomb explosion in Algeria resulted in 11 deaths and 31 injuries."
Causalities: explosion→death, explosion→injury
Example 2:
Text: "The earthquake killed 14 people, hundreds of people trapped in collapsed buildings."
Causalities: earthquake→killed, earthquake→collapsed, collapsed→trapped, earthquake→trapped
Example 3:
Text: "1. Warship INS Sukanya on Thursday foiled a piracy attempt in the Gulf of Aden between Somalia and Yemen. 2. The ship was escorting a group of five merchant vessels through the IRTC when the bid occurred."
Causalities: foiled→escorting, foiled→bid
In Example 1, the word "resulted" clearly indicates the causality between the events "explosion" and "death" as well as "injury," which is referred to as "explicit causality." In contrast, Example 2 lacks explicit connecting words or cues to indicate causalities, leading to three pairs of "implicit causality."
In the first two examples, all events are contained within the same sentence, termed "intra-sentence causality" or "sentence-level causality." In Example 3, the events "foiled," "escorting," and "bid" are not in the same sentence, creating what is known as "inter-sentence causality" or "document-level causality."
2.1.3 Event Causality Identification
Event causality identification (ECI) is a subtask within event relation extraction, aimed specifically at identifying the causalities between designated events. ECI focuses on extracting specific causal relationships, such as cause or caused_by. Based on the definitions of ERE in the literature [2], event causality identification can be formally defined as follows:
Definition 2: Event Causality Identification. Given a text $\mathcal{D}=[w_{1},\ldots,w_{n}]$ containing $n$ words and $k$ events $e_{1},e_{2},\ldots,e_{k}$ along with their mentions $m_{1},m_{2},\ldots,m_{k}$, where each event mention is a subset of $\mathcal{D}$. For any two events $e_{i}$ and $e_{j}$, determine their relationship type as one of {cause, caused_by, None}.
Based on whether the target event pairs are within the same sentence, ECI can be divided into two main research domains: Sentence-level Event Causality Identification (SECI) and Document-level Event Causality Identification (DECI).
2.2 Datasets
Annotated datasets are critical to advancing research in ECI. At present, several datasets are either specifically crafted for ECI tasks or can be adapted to evaluate ECI performance. Here, we provide a curated overview of 9 widely adopted datasets for ECI evaluation, detailing each dataset’s origin, version, and statistics.
  • SemEval-2010 Task 8 [35]: An enhanced version of SemEval-2007 [36] for classifying semantic relations between noun pairs, featuring nine relations (including causality) and an "OTHER" category. It emphasizes contextual clarity and reduces category overlap.
  • Chinese Event Causality (CEC) [37]: A Chinese causality dataset, extracted using layered Conditional Random Fields (CRFs). It includes many-to-many, explicit, inter-sentence, cross-sentence, cross-paragraph, embedded, and overlapping causalities.
  • Causal-TimeBank (CTB) [38]: Based on the TempEval-3 corpus [39], CTB tags events in TimeML (Temporal Markup Language) format. It uses $\mathtt{CLINK}$ tags for causal links between events and $\mathtt{C-SIGNAL}$ tags for words or phrases that indicate causality, with a focus on explicit causalities.
  • CaTeRS [40]: Based on 320 five-sentence stories from the ROCStories corpus [41], this dataset covers everyday scenarios with 13 types of causal and temporal relationships, focusing on narrative structure and story-driven causality.
  • BECauSE Corpus 2.0 [42]: An updated BECauSE corpus with texts from The Wall Street Journal, The New York Times, and the Penn Treebank. It includes detailed annotations of causal language and seven other semantic relations often linked with causality, focusing on explicit causalities.
  • EventStoryLine Corpus (ESL) [18]: Drawn from 22 topics in the ECB+ dataset on events like disasters, shootings, and trials, ESC supports cross-document reasoning and narrative generation. Available versions include v0.9, v1.0, v1.5, and v2.1.
  • FinReason [43]: A Chinese corpus in the financial domain aimed at extracting reasons for significant events in public company announcements. Each document may include multiple structured events, each with zero, one, or more reasons.
  • MAVEN-ERE [44]: The first large-scale, human-annotated ERE dataset, using documents from various topics on $\mathsf{Wikipedia}$. It builds on [45] with an enhanced annotation scheme to support various relationship types, including event coreference, temporal, causal, and sub-event relationships.
  • MECI [46]: The first multilingual dataset for event causality identification, sourced from Wikipedia in five languages (English, Danish, Spanish, Turkish, and Urdu). It offers valuable resources for multilingual recognition and cross-lingual transfer learning.
Table I summarizes the statistics of these datasets. In addition to the commonly used datasets mentioned above, there are also smaller, less frequently used, or non-open-source event causality datasets, such as Richer Event Descriptions [47], The Penn Discourse Treebank 2.0 [48], and CausalityEventExtraction  [^1] https://github.com/liuhuanyong/CausalityEventExtraction, which are not detailed here. To facilitate experimentation and comparison, we have organized these datasets and made them available on $\mathtt{Gitee}$ [^2] https://gitee.com/KGAWS/eci-related-works-and-datasets.
TABLE I: Statistics of Common ECI Datasets, "*" indicates Chinese datasets, "$\Delta$" indicates multi-lingual datasets.
Datase Document Count Event Count Relation Count Causality Count Year of Release
SemEval-2010 Task 8 10,717 9,994 10,717 1,331 2010
CEC* 332 5,954 340 340 2011
Causal-TimeBank 184 6,813 7,608 318 2014
CaTeRS 320 2,708 2,715 700 2016
BECauSE Corpus 2.0 121 2,386 1,803 1,803 2017
Event StoryLine (v0.9) 258 5,334 5,655 5,519 2017
FinReason* 8,794 12,861 11,006 11,006 2021
MAVEN-ERE 4,480 112,276 1,290,050 57,992 2022
MECIΔ 3,591 46,000 11,000 11,000 2022
2.3 Evaluation Metrics
As with most classification tasks, existing research commonly evaluates the performance of ECI using three metrics: precision, recall, and F1-score. Precision is defined as the percentage of true positives (causal event pairs) among all samples classified as positive by the model. Recall is calculated as the ratio of correctly identified positives to the total number of true positives. The F1-score, which is the harmonic mean of precision and recall, provides a balanced measure of precision and recall, evaluating the overall performance of causality identification.
Equations 13 present the formulas for these three evaluation metrics:
$\text{precision}=\frac{|\mathcal{P}^{t}|}{|\hat{\mathcal{P}|}},$ (1)
$\text{recall}=\frac{|\mathcal{P}^{t}|}{|\mathcal{P}|},$ (2)
[…]
where $\mathcal{P}^{t}$ represents the set of correctly identified causal event pairs, $\hat{\mathcal{P}}$ is the set of all identified causal event pairs, and $\mathcal{P}$ is the set of all true causal event pairs.
3 Key Technologies
This section covers four essential techniques for ECI: syntactic analysis, pattern matching, text embeddings (word-level and contextual), and graph embeddings. These techniques encompass all current methods for ECI, serving as the foundation for implementing and understanding related models. Figure 3 illustrates how these techniques are applied in ECI tasks.
Figure 3: Applications of key technologies in ECI tasks. Syntactic analysis and pattern matching can be directly combined for causality identification, while text and graph embeddings convert events into high-dimensional vectors, enabling deep learning algorithms to learn and infer causalities.
3.1 Syntactic Parsing
Syntactic parsing is an intuitive and straightforward approach for SECI. It analyzes a sentence’s grammatical structure by breaking it down according to grammar rules and building a syntactic tree or dependency graph. This process clarifies the structural relationships between elements like subjects, predicates, and objects. Syntactic parsing encompasses constituent parsing, dependency parsing, and semantic dependency analysis.
3.1.1 Constituent Parsing
Constituent parsing, based on Context-Free Grammar (CFG), breaks a sentence into phrases. CFG is defined as $G=(V,\Sigma,R,S)$, where $V$ is non-terminal symbols, $\Sigma$ is terminal symbols, $R$ is production rules, and $S$ is the start symbol. Parsing starts from $S$ and applies rules to form a parse tree, with the root as the sentence, leaves as words, and internal nodes as phrase components, see Figure 4.
Figure 4: Parse tree of the sentence “The food in the refrigerator spoiled because of the prolonged power outage, leaving a lingering unpleasant odor in the kitchen.” The parse tree indicates that the causal cue “because of” suggests a potential relationship between “outage” and “spoiled.” The red connections also imply a possible causal link between “odor” and “spoiled.”
3.1.2 Dependency Syntax Parsing
Dependency syntax parsing emphasizes grammatical relationships directly between words, without focusing on phrase structure. In this approach, a sentence is represented as a dependency tree, where each node is a word and edges illustrate dependencies between them. The root node typically represents the sentence’s predicate, with other nodes linked either to the root or to other words. Common dependency types include subject-verb, verb-object, and modifier relationships. Figure 5 shows an example of dependency parsing.
Figure 5: Dependency structure of the same sentence in Figure 4. Causality between "spoiled" and "outage" can be inferred from the path "spoiled-because of-outage" linked by the cue word "because of." Similarly, causality between "spoiled" and "odor" can be deduced from the path "spoiled-leaving-odor."
3.1.3 Semantic Dependency Parsing
Semantic Dependency Parsing (SDP) reveals both syntactic and semantic relationships in a sentence, producing a graph that captures the true meaning and logical links between words. It focuses on the semantic roles and event participation of words to uncover deeper meanings beyond surface syntax. Unlike traditional parsing methods that often use dynamic programming or graph search, semantic dependency analysis typically relies on deep learning models. Figure 6 shows an example of Abstract Meaning Representation (AMR) —— a technique of SDP.
Figure 6: A syntax graph generated using AMR. Based on this graph, the paths "spoiled-01 $\stackrel{{\scriptstyle arg0}}{{\longrightarrow}}$ outage" and "spoiled-01 $\stackrel{{\scriptstyle arg0}}{{\longleftarrow}}$ leaving-02 $\stackrel{{\scriptstyle arg1}}{{\longrightarrow}}$ odor" indicate potential causalities between "spoiled" and "outage" as well as between "spoiled" and "odor."
3.1.4 Summary
In traditional ECI models, syntax parsing is a key tool for understanding semantics. It effectively identifies word roles and relationships with minimal training, making it suitable for detecting clear, explicit causal links. With proper processing, it can also recognize cross-sentence causality. However, it struggles with implicit causality and may produce redundant links, and often requiring pattern-matching support.
3.2 Pattern Matching
Pattern matching uses predefined rules to identify segments that fit specific patterns, such as fixed strings, regular expressions, or syntactic structures. Patterns are defined based on the application needs, then matched with input text or syntax-parsed data. Regular expressions and part-of-speech (POS) tags are handled by annotation engines, while syntactic matching relies on graph-based similarity.
Pattern matching was a primary method in early text ECI tasks, effective for identifying explicit causal links but limited in detecting implicit ones. Its performance depends on syntax analysis results and is influenced by parser accuracy.
3.3 Text Embedding
Text embedding transforms language elements (words, sentences, or documents) into a continuous vector space, capturing semantic information for computational processing. With advancements in deep learning and PLMs, text embedding is essential for modeling complex semantics, identifying implicit causal links, and improving cross-domain generalization in ECI.
3.3.1 Word-Level Embedding
Word-level embeddings are essential for representing events in text, mapping each word to a vector. These embeddings fall into two types: static and dynamic.
Static embeddings (e.g., Word2Vec [49], GloVe [50]) assign fixed vectors to words, which limits their ability to handle polysemy and context sensitivity. They lack flexibility, as they can’t be fine-tuned for specific tasks.
Dynamic embeddings (e.g., BERT, RoBERTa [51]) are pre-trained on large corpora using self-supervised methods, capturing word dependencies and context through self-attention. These models generate context-aware vectors, handle polysemy, and can adapt to various downstream tasks through fine-tuning. They also create sentence-level embeddings to represent complex internal relationships.
Dynamic embeddings are generally applied to encode event mentions. These embeddings are then combined with the sentence’s global representation to create a context-aware event embedding. For an input sentence $X=[x_{1},...,x_{n}]$, the encoding process is as follows:
$\mathbf{X}=[\mathbf{x}_{0},\mathbf{x}_{1},...,\mathbf{x}_{n}]=\text{Encode}([x_{1},...,x_{n}]),$ (4)
where $\mathbf{x}_{i}$ is the embedding vector for each word $x_{i}$. For an event mention $e=[x_{m},...,x_{m+l}]$ (with $m>0$, $m+l\leq n$), the initial embedding $\mathbf{e}$ is calculated as:
$\mathbf{e}=\frac{1}{l+1}\sum_{i=1}^{l}\mathbf{x}_{m+i}.$ (5)
3.3.2 Contextual Embedding
Contextual embedding captures relevant contextual information, learning sentence-level features by filtering out unrelated details. In ECI tasks, it helps identify long-range and cross-sentence causalities. Typically, words are first encoded into vectors, then processed by sequence models to learn context-based features.
CNNs [52] capture local context through 1D convolutions with sliding windows, building from local to global context. RNNs [53] (and variants like LSTM [54], GRU [55]) process each word sequentially, carrying information forward to capture dependencies, while bidirectional RNNs [56] enhance context awareness with bidirectional modeling.
Transformer-based pre-trained language models (PLMs) use self-attention to process entire sequences, capturing long-range dependencies by adjusting word representations based on the full context. PLMs provide rich contextual embeddings, which can be fine-tuned to meet specific semantic requirements. In equation 4, $\mathbf{x}_{0}$ is the embedding of the [CLS] token (sentence embedding), which is also the contextual embedding.
3.3.3 Graph Embedding
In identifying causalities in long sentences and documents, irrelevant content can introduce noise, affecting performance. Modeling events as causality graphs addresses this issue. A causality graph is a directed graph represented as $\mathcal{G=(V,E)}$, where $\mathcal{V}$ consists of nodes (event mentions or other event-related features) and $\mathcal{E}$ consists of edges (causalities). Graph embedding techniques are increasingly used in ECI research. These techniques map nodes, edges, or entire subgraphs into a high-dimensional vector space, enabling embedded vectors to effectively capture relationships and features within the graph.
Graph embedding commonly uses Graph Neural Networks (GNNs) as encoders. GNNs operate on a message-passing mechanism, aggregating information from a node and its neighbors to update the node representation, reflecting both local structure and the node’s features. In a causality graph $\mathcal{G}$, for each node $v\in\mathcal{V}$, the neighboring nodes are denoted as $\mathcal{N}(v)$. The representation of node $v$ at layer $l$ is $\mathbf{h}_{v}^{(l)}$. The aggregation and update process in a GNN can be described as:
$\mathbf{h}_{v}^{(l+1)}=\text{UPDATE}^{(l)}(\mathbf{h}_{v}^{(l)},\text{AGGREGATE}^{(l)}(\{\mathbf{h}_{u}^{(l)}|u\in\mathcal{N}(v)\})),$ (6)
$\text{AGGREGATE}^{(l)}(\cdot)$ aggregates the representations of neighboring nodes at layer $l$, while $\text{UPDATE}^{(l)}(\cdot)$ updates the representation of node $v$ using its previous state $h_{v}^{(l)}$ and the aggregated neighbor information.
In ECI, commonly used GNNs include GCN [57], GAT [58], and GraphSAGE [59]. GCN captures graph structure by normalizing the graph Laplacian or adjacency matrix, aggregating neighboring nodes’ features to update the current node’s representation. GAT introduces a self-attention mechanism that assigns varying weights to neighboring nodes’ features, dynamically adjusting their influence on the target node based on importance. GraphSAGE is a scalable embedding model for large graphs. It samples neighboring nodes for embedding calculation, making it efficient for large-scale data.
4 Taxonomy
This section distinguishes SECI from DECI, presenting relevant methods and models. SECI and DECI are reviewed separately due to differences in task complexity, model demands, and challenges, with DECI being the more difficult. Methods are first categorized by their characteristics, followed by a discussion of their core techniques, classic models, strengths, and open challenges.
4.1 Sentence-Level Event Causality Identification
Sentence-Level event causality often conveyed through specific words and syntax, with many explicit cues and typical causal structures. Since SECI requires limited context, semantic interpretation is relatively straightforward. We categorize current SECI methods into four main types: feature pattern-based matching, deep semantic encoding, causal knowledge pretraining and prompt-based fine-tuning, and external knowledge enhancement.
4.1.1 Feature Pattern-Based Matching
Before deep learning became prevalent, feature pattern-based matching was the main approach for SECI. These methods leverage lexical cues, syntax, and pattern recognition to detect causal patterns and determine event relationships. This approach includes four types: template matching, syntactic patterns, co-occurrence, and feature-based classification. The general framework of feature pattern-based matching methods is shown in Figure 7.
Figure 7: The framework for feature pattern-based matching methods, where input sentences and events are preprocessed and parsed to generate semantic or syntactic structures and features, which then serve as the basis for feature pattern matching.
Template matching-based methods identify causalities using predefined patterns, such as role or graph templates (see Figure 8). [60] provides an in-depth analysis of causality templates and structures. [14] outlines five templates for causal patterns: (1) causal conjunctions, (2) causative verbs, (3) resultative complements, (4) conditional sentences, and (5) adverbs/adjectives with causal meaning, achieving a 68% identification rate on test data. [13] proposes constructing templates that represent roles and attributes within causal contexts. When text matches a template, matched words populate template slots to represent causalities. [61] enhances accuracy and scalability by using syntactic graph templates, which align with parse trees to pinpoint cause and effect segments in text, making template matching more precise. Template matching methods are interpretable and sensitive to explicit causalities. However, they can only identify causalities that fit predefined templates, making it challenging to address implicit, complex, or irregular expressions of causality.
Figure 8: The general framework of template matching methods. Dependency graph or syntax graph/tree of the sentence is generated using a parser, which is then matched against predefined templates to identify causalities.
Syntactic pattern-based methods identify causalities using predefined structures. Ishii and Ma [62, 63] created a topic-event causal model that represents text as a causal network graph. Their incremental approach merges similar event vertices based on thematic and event keywords, calculating importance scores to optimize the network. Caselli and Vossen [18] introduced the Order Presentation (OP) method, pairing each event with subsequent events based on their order, assuming earlier events may cause later ones (also for DECI). Girju and Moldvan [64] developed a semi-automated method to identify causal syntactic patterns. They extract noun phrase pairs from $\mathsf{WordNet}$ [65] and search for sentences containing these pairs. They then identify sentences that fit the pattern "NP1 verb/verb expression NP2", assessing whether the pattern indicates causality. Ittoo and Bouma [66] modeled events as a causality graph, using shortest paths to represent relationships corresponding to syntactic patterns, then iteratively identified linking patterns and scored their reliability based on event pair purity. Building on this, Ittoo et al. [67] proposed a selection strategy for causal event pairs to ensure diverse representation of causality subtypes. Syntactic pattern-based methods can identify more complex causalities than template matching. They automatically generate patterns using parsing tools, minimizing manual construction. They can also differentiate causal meanings and resolve ambiguities. However, reliance on parsers for analysis increases computational demands and may affect performance if the parser is inaccurate.
Co-occurrence pattern-based methods identify causalities based on event co-occurrence frequency and semantic associations. Do et al. [19] proposed a minimally supervised learning approach that uses pointwise mutual information and inverse document frequency to measure relationships between events. They calculate co-occurrence counts to predict causalities and leverage discourse connectives for additional context. They then use Integer Linear Programming (ILP) to ensure causal consistency. Caselli and Vossen [18] employed Positive Pointwise Mutual Information (PPMI) to evaluate the semantic relevance of event pairs by analyzing co-occurrence frequencies in large datasets. The improved PPMI-CONTAINS model adds temporal constraints, requiring event pairs to have PPMI values within a specified range and to share the same time frame. Co-occurrence-based methods use statistical information from large corpora instead of predefined templates, enabling them to address implicit causalities and adapt to different domains. However, reliance on statistical associations can result in many false positives.
Feature classification-based methods utilize machine learning techniques to model ECI as a binary classification task. Zhao et al. [68] introduced method to measure similarity in sentences with causal connectives and categorizing these connectives to enhance understanding of causality. They ultimately used a naive Bayes model for causal learning. Mirza and Tonelli [17] inferred causal chains by analyzing causal verbs and their dependencies, using feature-based classification to identify causalities. They also proposed an SVM-based model [69], which uses lexical-syntactic features and dependency paths for classification. Mirza [16] incorporated temporal rules and introduced causal signals and causal links between events, utilizing features from lexical-syntactic information and context for classification. Aforementioned models automatically learn diverse feature combinations to improve accuracy. However, their performance declines significantly when annotated data is sparse. Furthermore, these models often capture only shallow semantic features.
Open Challenges. Pattern matching-based methods are simple, efficient, and interpretable, with low computational demands, making them popular in early research. However, they lack the ability to capture deep semantic information and rely heavily on external parsers and predefined lexical or syntactic patterns. This dependence limits their scalability, flexibility, and generalizability, causing performance to degrade significantly with complex semantics, unfamiliar structures, or noisy text.
4.1.2 Deep Semantic Encoding
With advancements in deep learning, neural sequence models have enabled richer representations for text and events, particularly in managing long-distance dependencies. PLMs generate context-sensitive, dynamic word embeddings through large-scale pre-training, thereby capturing more intricate semantic relationships.
Deep semantic encoding-based ECI methods employ deep sequence models or PLMs as encoders, applying tailored encoding strategies to capture the semantic and contextual information of events and generate high-dimensional embeddings. The embeddings of two target events are concatenated and input into a classifier (e.g., a multilayer perceptron), which learns the features of these high-dimensional embeddings through neural networks to assess the presence of causality (see Figure 9).
Figure 9: Framework of deep semantic encoding for ECI. Text and event mentions, along with additional features (such as semantic structure, event information, and labels), are encoded as embeddings and fused. A classifier is then used to generate the predicted probability for the event relation type.
4.1.2.1 Simple Encoding Strategy
Simple encoding strategies leverage sequence models (e.g., RNNs, LSTMs) and PLMs (e.g., BERT, RoBERTa, and Longformer [70]) as encoders to obtain contextual embeddings of event pairs, which are then input into a classifier to determine their relationship type. Although few of them are specifically covered in dedicated literature [71], these methods are frequently used as baseline comparisons in research.
For sequence models, an input sentence $X=[x_{1},\dots,x_{n}]$ is processed by first obtaining static word embeddings for each word, followed by encoding with a sequence model to capture contextual information, generating event embeddings as follows:
$\mathbf{x}_{i}=\mathrm{Emb}(x_{i}),$ (7)
$\widetilde{\mathbf{X}}=\left[{\widetilde{\mathbf{x}}}_{1},{\widetilde{\mathbf{x}}}_{2},\ldots,{\widetilde{\mathbf{x}}}_{n}\right]=\mathrm{SeqModel}\left(\left[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\right]\right),$ (8)
$\mathrm{Emb}$ and $\mathrm{SeqModel}$ denote static embedding model encoding and sequence model encoding, respectively. The event embeddings are calculated based on the words corresponding to event mentions. For PLMs, event embeddings are derived using equation 4.
The embeddings of two events, $\mathbf{e}_{i}$ and $\mathbf{e}_{j}$, are concatenated to form the event pair representation, either as $\mathbf{ep}_{ij}=[\mathbf{e}_{i};\mathbf{e}_{j}]$ or $\mathbf{ep}_{ij}=[\mathbf{e}_{\text{[CLS]}};\mathbf{e}_{i};\mathbf{e}_{j}]$. This concatenated vector $\mathbf{ep}_{ij}$ is then input to a classifier, typically a Multi-Layer Perceptron (MLP), as follows:
$\mathbf{p}=\mathrm{MLP}(\mathbf{ep}_{ij})=\mathrm{softmax}(\mathbf{W\cdot ep}_{ij}+\mathbf{b}),$ (9)
$\mathbf{W}$ and $\mathbf{b}$ are learnable parameter matrices, and $\mathbf{p}\in\mathbb{R}^{2}$ (ignoring direction) or $\mathbf{p}\in\mathbb{R}^{3}$ (considering direction). Each dimension of $\mathbf{p}$ corresponds to the probability of the causality or non-causality.
During training, parameters are optimized on a annotated dataset using cross-entropy loss to enable causality classification:
$\mathcal{L}=-\sum_{j=1}^{N}\sum_{i=1}^{C}y_{i}^{j}\log\left(p_{i}^{j}\right),$ (10)
$C=2$ or $3$ corresponds to the output dimension of $\mathbf{p}$, $y_{i}^{j}$ represents the true label, $p_{i}^{j}$ the predicted probability, and $N$ the number of labeled event pairs. Encoder parameters can also be updated as needed.
Simple encoding strategies have limited capacity for capturing context and often generate low-quality embeddings. They require large amounts of labeled data for training, are prone to false positives, and exhibit relatively weak cross-domain generalization.
4.1.2.2 Enhanced Encoding Strategy
Enhanced encoding strategies leverage sequence models or PLMs combined with rich feature to capture context, improving the quality of event and event-pair representations. These strategies integrate additional information into embeddings before classification. Current research include three main approaches: semantic structure enhancement, event information enhancement, and causal label enhancement. Table II summarizes the enhancement details of these encoding models.
Semantic structure enhancement encoding strategies use structured semantic information, like dependency relations, syntactic paths, and semantic graphs, to enrich event representations. Tri-CNN [72] combines syntactic and contextual information through a three-channel CNN, integrating paths and event-pair embeddings for a comprehensive global representation. SemSIn [24] captures implicit relationships by encoding events with BERT, parsing text into semantic graphs with AMR, and aggregating event and neighbor information using GNNs. SemDI [73] uses sentence-level semantic dependencies with three modules: RoBERTa for word dependencies, a Cloze test for context prediction, and multi-head attention to evaluate causal associations. These strategies, although effective for handling complex sentences, can be computationally demanding, particularly when dealing with long or cross-sentence causalities. This results in increased training and inference times.
Event information enhancement encoding strategies improve event representations by capturing various elements such as semantics, location, time, and participants, enriching the context and detail of each event. Kadowaki et al. [23] use BERT to encode event-specific elements (e.g., roles, positions) and train multiple classifiers to learn different annotator perspectives, combining their outputs for final predictions. DualCor [74] propose a dual grid tagging scheme with BERT to capture event correlations, integrating event type information with contextual representations to distinguish cause-effect pairs. Yang et al. [75] employ a Bi-LSTM-based Siamese network to incorporate event triggers and associated arguments (like time and location), enhancing event embeddings. A knowledge graph-inspired approach is then used to infer causalities based on event similarity. These methods incorporate multiple event elements and features, enriching the representation of events and causalities. However, they are computationally intensive and require well-designed event elements and features to be effective.
Label information enhancement encoding strategies refine event pair embeddings by using causal labels, applying techniques like contrastive learning, graph interactions, and counterfactual reasoning. SCL [76] enhances event representations by leveraging causal labels, constructing triplets (positive and negative event pairs) and uses contrastive loss to create more distinctive embeddings. ECLEP [77] combines causal label information with event pair interactions. ECLEP builds an event-pair interaction graph, integrating causal labels and using GAT to update representations. CF-ECI [78] uses counterfactual reasoning to mitigate bias in ECI. The model assesses biases from context and event pairs separately, using causal labels to adjust predictions through weighted debiasing. These methods maintain high accuracy even with noisy data or ambiguous context. However, they rely heavily on causal labels and require extensive labeled data.
TABLE II: Enhanced Features of Different Enhanced Encoding-based Models
Enhance Strategy Model Encoder Enhance Features
Semantic structure Tri-CNN [72] RCNN Syntactic information, contextual information
SemSIn [24] BERT/GNN/BiLSTM Event-centric structure, event association structure
SemDI [73] RoBERTa Semantic dependency
Event information BERT+MA [23] BERT Event location, text segment, features (roles)
DualCor [74] BERT Event type and arguments
Siamese BiLSTM [75] BiLSTM Event time, location, participants
Label information SCL [76] BERT Labeled positive and negative samples
ECLEP [77] BERT/GAT Causally related event annotations
CF-ECI [78] BERT Causal labels
Summary. Simple encoding strategies are easy to implement and train, offering high computational efficiency but are only suited for clear, low-noise texts within a single domain. Enhanced encoding strategies provide a more comprehensive modeling of event causality, improving accuracy and robustness in ECI, especially in complex contexts and long-distance dependencies. They are effective in handling complex semantic structures and exhibit stronger cross-domain generalizability. However, deep semantic encoding-based methods require a substantial amount of labeled data. Without sufficient supervised learning, these models struggle to learn high-quality encoding and classification capabilities, which impacts inference accuracy.
4.1.3 Causal Knowledge Pretraining and Prompt-Based Fine-Tuning
With advancements in PLMs and prompt learning, two new ECI approaches have emerged: (1) causal knowledge pre-training, which involves retraining or continually pretraining PLMs with causal commonsense or domain-specific knowledge, enabling zero-shot, end-to-end ECI using simple prompts; and (2) prompt-based fine-tuning, which uses standard PLMs with tailored prompt templates and prompt learning algorithms, trained on labeled datasets, to guide end-to-end ECI with enhanced accuracy through dynamic prompts.
4.1.3.1 Causal Knowledge Pre-training
Causal Knowledge PLMs (CKPLMs) enhance causal reasoning through extensive causal knowledge bases and cross-task training. These models can be applied to ECI tasks with simple prompts, requiring no fine-tuning, and exhibit strong cross-domain generalizability.
CausalBERT [79] collects multi-level causal pairs from unstructured text and injects this knowledge into BERT to improve causal reasoning. Its trained in three stages: (1) unsupervised pre-training with a BERT-based masked language model; (2) causal pair classification and ranking with data from sources like $\mathsf{CausalBank}$ [^3] https://github.com/eecrazy/CausalBank/ and $\mathsf{ConceptNet}$[80]; and (3) application to causal classification tasks using simple prompts or fine-tuning for causal reasoning tasks.
Causal-BERT [81] enhances implicit causality detection by using sentence and event contexts. It consists of three frameworks: (1) C-BERT, which encodes events and sentence contexts to detect causality; (2) event-aware C-BERT, which integrates event and sentence contexts for improved predictions; and (3) masked event C-BERT, which replaces events with “BLANK” to learn implicit structures, later fine-tuned with event-aware C-BERT.
UnifiedQA [82] is a multi-task QA model trained on various question formats, demonstrating strong cross-domain performance. The latest version, UnifiedQA-v2, expanded its training data to include 20 datasets with causal knowledge, further improving its generalizability.
LLMs also function as CKPLMs due to their extensive training on diverse corpora rich in causal information. Gao et al. [27] designed Chain of Thought (CoT) and In-Context Learning (ICL) prompts to guide ChatGPT for ECI. CoT prompts encourage step-by-step reasoning, while ICL provides examples to shape output. However, LLMs are prone to “causal hallucination,” often over-attributing causality, and their ECI performance is highly sensitive to prompt design, which can lead to variable outcomes.
Summary. CKPLMs enable rapid ECI with minimal labeled data, effectively handling implicit causality and generalizing well across domains. However, challenges like causal hallucination and prompt sensitivity lead to high false positive rates and potential misinformation. When the knowledge in the text conflicts with the knowledge in CKPLMs, it may result in incorrect responses. Although LLMs have improved zero-shot performance in event-related tasks, achieving high-accuracy ECI still requires carefully designed algorithms to capture complex causal patterns.
4.1.3.2 Prompt-Based Fine-Tuning
Prompt-based fine-tuning methods utilize the semantic understanding and extensive knowledge of PLMs by employing fixed or dynamic prompt templates. By framing ECI as a masked prediction or multiple-choice task, these methods enable PLMs to directly generate causal labels. Table III shows the prompt templates of these methods.
TABLE III: Prompt templates and PLMs of prompt-based fine-tuning for ECI.
Model Prompt Template Notations
MRPO [83] From context [D], it is [MASK1] that E1 has a causal relationship to E2. [Es][Et][/Es][/Et]: boundary markers. $\mathcal{T}_{a}(d_{i})$: the i-th demonstration example template. $\mathcal{T}_{p}(q)$: prompt template for the query event pair. $[MASK1]$in {Yes,No} $[MASK2]$in {Cause,Causedby,None} $[MASK3]$: causal cue words $[MASK4]$: event mention
KEPT [84] [Es] E1 [/Es] [MASK2] [Et] E2 [/Et].
DPJL [85] (1) In this sentence, E1 [MASK2] E2. (2) The cue word of E1 [MASK2] E2 is [MASK3]. (3) According to the sentence,E1 causes [MASK4],E1 is caused by [MASK4].
ICCL [86] [CLS] $\mathcal{T}_{a}(d_{1})$ [SEP] … $\mathcal{T}_{a}(d_{n})$ [SEP] $\mathcal{T}_{p}(q)$ [SEP]
GenECI [87] [D] Is there a causal relation between E1 and E2?
GenSORL [88] [D] Is there a causal relation between [Es] E1 [/Es] and [Et] E2 [/Et]?
RE-CECI [89] In this sentence, E1 [MASK2] E2.
MRPO [83] converts ECI into a masked language modeling task. A predefined template with a [MASK] token is used for causal classification. BERT predicts the mask, and an MLP processes this output to classify event causality, considering all potential values for the mask, which reduces reliance on specific keywords. KEPT [84] incorporates external knowledge from $\mathsf{ConceptNet}$ to enhance ECI, using a prompt template that highlights events with boundary tokens. BERT encodes this structure, and selective attention combines relational information from $\mathsf{ConceptNet}$ with event representations. DPJL [85] frames ECI as a mask prediction task with templates. DPJL adds two auxiliary prompts for cue word detection and causal event detection, respectively. Joint training on these two tasks is then leveraged for model learning. ICCL [86] combines prompt-based learning with in-context contrastive learning. The model uses RoBERTa to predict [MASK] for the query template. The contrastive learning module retrieves positive and negative samples, refining the model’s ability to differentiate causal from non-causal pairs. GenECI [87] uses a generative approach with T5 [90] to predict causal labels and generate context words connecting events in the template, generating causal paths as the output. Reinforcement learning optimizes generated output by rewarding accuracy, similarity, and alignment with input structure. GenSORL [88] is an improvement on GenECI, which extracts simplified dependency paths from syntax trees to reduce noise in causality prompts. It applies a policy gradient with baseline for more stable training and includes an auxiliary reward to improve similarity between generated outputs and input sentences. RE-CECI [89] implements a retrieval-enhanced framework for Chinese ECI by retrieving examples from a stored database using k-nearest neighbors. These retrieved examples are incorporated into a prompt template. RoBERTa then encodes this prompt structure to predict causality.
Summary. Prompt-based fine-tuning leverages PLMs’ context understanding and reasoning abilities, adapting well to low-data scenarios, and flexible prompt designs improve cross-domain generalization. However, models may over-rely on language patterns, which can lead to false negatives for implicit causality. Prompt quality is crucial, as complex contexts or noise can impact robustness if prompts are not well-designed.
4.1.3.3 Open Challenges
Causal knowledge pre-training and prompt-based fine-tuning methods reduce the need for extensive labeled data and show strong cross-task transferability. However, they face challenges such as causal hallucination, prompt dependency, and knowledge conflicts. This indicates that, while these PLM-based methods improve efficiency in ECI, further refinement in algorithm and prompt design is needed to achieve higher accuracy and robustness.
4.1.4 External Knowledge Enhancement
ECI models that rely solely on event text require extensive labeled datasets, which are challenging to obtain. Causal expressions also vary across topics and contexts, limiting model performance and increasing overfitting risk. To address this, recent studies have integrated external knowledge, primarily through Knowledge Graphs (KGs), which provide causal background knowledge to alleviate data scarcity and support reasoning (see Figure 10). External knowledge-enhanced methods can be categorized into data augmentation, representation enhancement, and reasoning enhancement.
Figure 10: Framework of external knowledge enhancement for ECI. KGs generate supplementary annotated data enhance model training. Events in KGs are encoded via graph embeddings and merged with contextual embeddings of event mentions, improving representation quality. Additionally, KG subgraphs and reasoning paths support causal reasoning.
4.1.4.1 Data Augmentation
Data augmentation with external knowledge addresses limited training data by generating samples with similar semantic patterns to the target dataset, enriching training data for causal learning and causality identification.
KnowDis [91] uses distant supervision to expand causal event pairs with $\mathsf{WordNet}$ and $\mathsf{VerbNet}$ [92], applying a commonsense filter to remove noise and a re-annotation strategy that gradually incorporates noisy data through self-training and annealing. LearnDA [93] combines sentence generation and causality identification through dual learning. It generates sentences based on causalities from $\mathsf{WordNet}$ and $\mathsf{VerbNet}$, and then maps sentences back to identify causality. This dual learning approach creates new, relevant data to further train causality models. CauSeRL [94] employs self-supervised learning to capture context-specific causal patterns from external knowledge bases (e.g., $\mathsf{ATOMIC}$ [95]). Using a contrastive transfer approach, it aligns learned patterns from a teacher model with a student model for ECI, maximizing mutual information to transfer causal knowledge effectively.
Summary. Data augmentation enriches causal samples, improves understanding of complex causal patterns, reduces overfitting, and decreases dependence on labeled data. However, its effectiveness depends on the knowledge base’s coverage and quality; commonsense knowledge alone may be insufficient for domain-specific ECI. Excessive augmentation may also introduce noise, affecting model performance.
4.1.4.2 Representation Enhancement
Representation enhancement methods enrich event pair representations by integrating external knowledge and contextual information. These methods typically use PLMs or GNNs to incorporate knowledge base information into event representations.
KnowMMR [25] combines external knowledge and event masking to improve generalization in ECI tasks. It retrieves relevant knowledge from $\mathsf{ConceptNet}$, encodes it with BERT to enhance event pair embeddings, and applies event masking to learn context-independent patterns, allowing the model to better handle new or unseen causality patterns. ECIFF [96] enhances event representation by combining contextual, semantic, and syntactic features. It encodes causal triples from $\mathsf{CauseNet}$ [97] and uses adversarial learning to create event pair embeddings, incorporating both text and features of KG for a more robust event representation. DSET [98] uses dual semantic enhancement with temporal prompts. It integrates temporal cues and external semantic information from $\mathsf{ConceptNet}$ to enhance event representations, using RoBERTa to encode relevant knowledge and GraphSAGE to capture dependency structures in event graphs. DFP [99] fuses commonsense and event semantics through a Distill, Fuse, and Pre-train approach. It applies GAT to process KG information and BERT for event-related text, continuously pretraining the model on MLM, concept triple completion, and text-graph contrastive learning tasks to improve representation. ECIHS [100] proposed a model to capture syntactic, semantic, and background knowledge by encoding text with Transformers, dependency structures with syntax trees, semantic information with AMR graphs, and background knowledge from $\mathsf{ConceptNet}$. These layers are fused using Graph Transformer [101] and GCN to create enhanced event representations. KLPWE [102] combines external knowledge and multi-level word embeddings to enhance individual event representations. It integrates KG embeddings with morphological and POS features, as well as BERT-based dynamic embeddings, to predict causal roles for each event.
Summary. Compared to data augmentation, representation enhancement focuses on improving model generalization and adaptability to new tasks. However, its effectiveness depends significantly on the quality of external knowledge and the integration approach; poorly integrated knowledge can introduce noise, hindering inference accuracy.
4.1.4.3 Reasoning Enhancement
Reasoning enhancement methods leverage external knowledge to construct relational graphs or causal chains between events, facilitating the reasoning of implicit causalities. These approaches typically use multi-hop or logical reasoning mechanisms to establish deeper causal connections.
LSIN [103] enhances causal reasoning by integrating knowledge from $\mathsf{ConceptNet}$ and $\mathsf{COMET}$ [104], constructing descriptive graphs with one-hop neighbors and multi-hop relational graphs between events. Variational inference is used to learn latent inference structures, which are combined with BERT-encoded event representations. Chen and Mao [105] propose a model that combines explicit knowledge from $\mathsf{FrameNet}$ [106] and implicit knowledge from $\mathsf{COMET}$ to improve complex causality understanding. It builds an event pair graph with GCN for event interactions and uses causality-related words and inferred causal facts for enhanced reasoning. C3Net [107] incorporates intuitive and logical reasoning by retrieving causal paths from external knowledge, processing them through association, review, and inference, and applying a competitive mechanism to select the most reliable inference. GCKAN [108] uses external knowledge to build event neighbor and path graphs, with GAT aggregating information from $\mathsf{ConceptNet}$. Contrastive learning is utilized to reduce noise, strengthening the robustness of event representations.
Summary. Reasoning enhancement improves cross-domain causal understanding by establishing explicit and implicit links between events. However, its effectiveness depends on the coverage and reliability of the knowledge base; excessive reliance on external knowledge may limit the model’s ability to manage ambiguous or conflicting causal information. In addition, the reliability of the reasoning path is often hard to ensure, and there may be error propagation.
4.1.4.4 Open Challenges
External knowledge enriches context, mitigates causal data scarcity and reduces overfitting, improving model generalization, especially in cross-domain tasks. However, model performance relies on the knowledge base’s coverage and quality; insufficient or conflicting knowledge can mislead causal inference. Integrating external knowledge may also introduce noise or redundancy, increasing complexity. In inference enhancement, incorrect multi-hop paths can propagate errors, affecting accuracy.
4.2 Document-Level Event Causality Identification
As ECI tasks become more complex, DECI has emerged as a primary research focus. Document-level causal links often span multiple sentences or paragraphs, requiring models to integrate broader context, resolve coreferences, and handle implicit causality. This complexity limits the effectiveness of traditional syntax and pattern-based methods, necessitating advanced semantic understanding and reasoning. Current DECI methods primarily include event graph-based models and prompt-based based models, with some models also using knowledge enhancement, text encoding, and data generation techniques.
4.2.1 Event Graph Reasoning-Based Methods
Event graph reasoning-based DECI methods model document-level causalities by constructing diverse event-related graph structures, framing causality as a reasoning task between nodes and edges (see Figure 11). Typically, these methods first use PLMs to encode events and context, then build a heterogeneous graph centered on events, treating elements like events, words, and phrases as nodes. Various types of edges (e.g., event co-occurrence, syntactic dependencies, and semantic similarities) capture relationships between nodes. Graph embedding models then aggregate node information, updating event representations to capture global context and infer causalities through high-order or path-based reasoning.
Figure 11: Framework of event graph reasoning for DECI. Event graphs can be homogeneous or heterogeneous, representing various information types—event mentions, event pairs, documents, sentences, and words—as nodes connected by different predefined edges (dark lines, light lines, and dashed line). Graph embedding models encode these nodes, applying diverse strategies to create embeddings for events and event pairs, which are then used by a classifier to predict event causalities.
DCGIM [109] models events as nodes and edges as causal links between them. To capture document-level context, the model uses attention mechanisms to update sentence representations and integrates the contextual information of each event. A directed GCN aggregates and updates node information, and a biaffine layer predicts edge types. RichGCN [110] constructs a document-level interaction graph with nodes representing words, events, and entities, and edges capturing discourse relations, syntactic dependencies, and semantic similarity. A GCN aggregates information along these edges, using sparse graph regularization to reduce noise, and an MLP classifies causalities. GESI [111] builds a heterogeneous graph with nodes for event mentions, events, and documents, and edges for mention-mention, mention-event, event-event, and event-document connections. GCNs aggregate node information, while attention mechanisms weight mentions of the same event. The model performs path-based reasoning on event subgraphs to infer causalities. ERGO [112] creates a high-order event relational graph, where nodes are event pairs and edges represent potential causal links. It uses a Graph Transformer with self-attention to capture high-order interactions and applies cross-graph constraints to reduce false positives, while an adaptive focal loss reduces false negatives in challenging samples. CHEER [113] builds a high-order event interaction graph with nodes for events and event pairs, using edges to capture event-event and event pair interactions. The model leverages centrality information to identify key events, applies multi-layer GCNs to aggregate information, and uses self-attention to dynamically adjust edge weights. DocECI [114] constructs two graphs: a text structure graph with nodes for words and sentences, and a mention relation graph with nodes for event mentions and sentences. Edges capture word adjacency, dependency, sentence co-occurrence, and coreference. A Relational GCN aggregates information across the graphs, using multi-head attention to integrate causal information across subspaces. PPAT [115] builds a sentence-boundary event relationship graph with nodes representing event pairs and edges capturing shared events. The model performs local context-based reasoning for intra-sentence causalities and infers inter-sentence causalities progressively, using a pairwise attention mechanism to focus on relevant causal chains. iLIF [116] constructs an event causal graph with nodes for events and edges for intra- and inter-sentence links. Through iterative learning, it updates event representations based on high-confidence causalities, ensuring the model focuses on reliable intra-sentence causal links, which then inform inter-sentence reasoning. GIMC [117] is designed for zero-shot cross-lingual DECI, building a heterogeneous graph with nodes for event pairs, sentences, phrases, and statements, and edges for connections such as phrase-phrase, sentence-phrase, and event-event links. A GAT models interactions, with multi-granularity contrastive learning aligning causal representations across languages at the statement and event levels. We present the types of nodes and edges used in the graphs of the above models in Tabel IV.
TABLE IV: The node and edge types in event graph reasoning-based DECI models.
Model Nodes Edges
DCGIM [109] Events, Context Causalities
RichGCN [110] Words, Event Mentions, Entity Mentions Discourse Association, Syntactic Dependency, Semantic Similarity
GESI [111] Event Mentions, Events, Document Mention-Mention, Mention-Event, Event-Event, Event-Document
ERGO [112] Event Pairs Causalities
CHEER [113] Events, Event Pairs Event Pair-Event Pair, Event-Event, Event Pair-Event
DocECI [114] Words, Sentences, Event Mentions Word Adjacency, Word Dependency, Word-Sentence, Sentence Adjacency, Sentence Supplement
PPAT [115] Event Pairs Shared Event
iLIF [116] Events Intra-sentence, Inter-sentence
GIMC [117] Event Pairs, Sentences, Phrases, Utterances Phrase-Phrase, Sentence-Phrase, Phrase-Event, Sentence-Event, Utterance-Event, Event-Event
Summary. Event graph-based ECI methods effectively model long-distance dependencies by aggregating information across sentences and paragraphs, updating event representations with a global perspective and reducing noise impact on causality identification. However, as document size grows with more events, phrases, and elements, the graph size increases, potentially introducing unnecessary edges that add noise and affect model accuracy. Additionally, diverse node and edge types require specialized aggregation and reasoning mechanisms, complicating model structure and making practical tuning more challenging.
4.2.2 Prompt-based methods
Seem to models for SECI, prompt-based methods DECI methods use PLMs to convert ECI tasks into generation or mask-filling tasks through carefully designed natural language prompts. By providing clearly defined prompts, these methods dynamically capture context, guiding the model to predict or generate causal links between events.
CPATT [118] treats ECI as a natural language generation task. By applying a constrained prefix attention mechanism, it reduces co-occurrence interference and prompt ambiguity. Specific templates for events and relations are used to dynamically generate context-relevant prompts, enhancing the capture of semantic features across relationship types. A scoring algorithm interprets the template outputs to predict causalities. DAPrompt [119] bypasses the limitations of manually designed templates by assuming a causal link between events, then testing this through plausibility evaluation. The prompt includes event placeholders and causal assessment masks, with two MLM classifiers calculating the probability of each event matching its placeholder. Causality is determined based on these probabilities and a preset threshold. HOTECI [120] leverages optimal transport at sentence and word levels to select key context relevant to causality. Using a prompt to ask if a causality exists, the model outputs essential contextual words and a causality label (“yes” or “no”). Reinforcement learning further optimizes context selection by rewarding choices that enhance causal identification accuracy.
Summary. Prompt-based methods leverage extensive knowledge in PLMs and flexible prompt design to adapt to diverse expressions of event relations, offering strong generalization. However, compared to event graph-based methods, prompt-based approaches are better suited for short-distance causal links. When causalities span multiple sentences or paragraphs, these models may struggle with long-distance dependencies due to noise interference.
4.2.3 Other Methods
In addition to the main approaches, some studies use strategies such as linear programming, external knowledge enhancement, sparse syntactic attention, and diffusion models to tackle DECI tasks.
Cheng and Miyao [121] apply dependency tree structures to represent event relations, linking cross-sentence events through a “common root” shared by adjacent sentences. A BiLSTM encoder processes each dependency path to capture relationship features and predict causality. Gao et al. [122] integrate document-level structures by identifying “main events” likely involved in multiple causal links, combined with syntactic, discourse, and coreference constraints. An ILP framework optimizes causal link predictions across event pairs. KADE [123] enhances event representations by retrieving related knowledge from $\mathsf{ConceptNet}$ and merging it with the original event sentence. For rare events, KADE introduces an analogy mechanism, using k-Nearest-Neighbors (kNN) to retrieve similar events from memory and refine representation. SENDIR [124] applies sparse attention to selectively focus on relevant information within long texts, differentiating intra- and inter-sentence causality. SENDIR builds reasoning chains for cross-sentence events by identifying supporting events through gated attention. DiffusECI [125] iteratively adds and removing noise to transform event context representations into refined causal labels. During inference, the model progressively filters noise to enhance causal signal clarity, improving causality predictions across lengthy texts.
DECI research is still in its early stages, and the methods discussed above are not yet part of a cohesive methodological framework. Key challenges in DECI include causal generation mechanisms, noise reduction in long texts, and high-quality event representations. As these models evolve, future research will likely focus on refining these aspects to improve DECI performance.
4.3 Performance Assessment
We performed a comprehensive analysis and discussion of all the aforementioned ECI models, comparing them across five key aspects: effectiveness in identifying implicit causality, data annotation requirements, cross-domain generalizability, reliance on external knowledge/tools, and issues related to false positives. The results of this comparison are summarized in Table V. Note that the results presented in the table are derived from a qualitative, relative comparison, based on an analysis of both the model frameworks and their experimental results. These results are intended to serve as a reference only.
TABLE V: Evaluation includes five evaluation aspects and all the results are relative. "I" assesses the ability of a method to identify implicit causalities($\star$: Strong ability $\sqrt{}$: Moderate ability ×: Weak ability). "A" indicates the level of annotated data required for effective performance (×: High requirement $\sqrt{}$: Moderate requirement $\star$: Low requirement). "G" measures the method’s ability to generalize across different topics or domains ($\star$: Strong generalization capability $\sqrt{}$: Moderate capability ×: Weak capability). "E" indicates the degree to which a method relies on external knowledge (↑: High dependency →: Moderate dependency -: No dependency). "F" assesses the extent to which a method incorrectly reports false positives (×: Severe false $\sqrt{}$: Minimal false positives $\star$: Little to no false positives).
Task Taxonomy Model Conference/Journal Year I A G E F
SECI Feature Pattern -Based Matching Template Matching - - × $\star$ × - $\star$
Syntactic Pattern - - × $\star$ × - $\star$
Co-occurrence Pattern - - $\sqrt{}$
Feature Classification - - $\sqrt{}$
Deep Semantic Encoding Enhanced Encoding Tri-CNN [72] ACL 2016 $\sqrt{}$ $\sqrt{}$
BiLSTM [56] EMNLP 2017 $\sqrt{}$
Siamese-BiLSTM [75] ICONIP 2018 $\sqrt{}$ $\sqrt{}$
BERT+MA [23] EMNLP 2019 $\sqrt{}$ $\sqrt{}$
SemSIn [24] ACL 2023 $\star$ $\sqrt{}$star - $\sqrt{}$
SCL [76] CCIS 2023 $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$
ECLEP [77] ACL 2023 $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$
CF-ECI [78] ACL 2023 $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$
SemDI [73] arXiv Preprint 2024 $\star$ $\sqrt{}$star - $\star$
Causal Knowledge Pre-training and Prompt-based Fine-tuning Causal Knowledge PLM Unified-QA [82] EMNLP 2020 $\sqrt{}$star $\star$ - ×
CausalBERT [79] arXiv Preprint 2021 $\sqrt{}$star $\star$ - ×
Causal-BERT [81] arXiv Preprint 2021 $\sqrt{}$star $\star$ - ×
LLM-ECI [27] EMNLP 2023 $\star$ $\star$ $\star$ - ×
Prompt-based Fine-Tuning for PLM DPJL [85] COLING 2022 $\star$ $\sqrt{}$star - $\star$
GENECI [87] ACL-SIGLEX 2022 $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$
KEPT [84] KBS 2023 $\sqrt{}$star $\star$ $\sqrt{}$
MRPO [83] IJCNN 2023 $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$
RE-CECI [89] NLPCC 2023 $\sqrt{}$star $\sqrt{}$ $\sqrt{}$
GenSORL [88] KBS 2024 $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$
ICCL [86] arXiv Preprint 2024 $\sqrt{}$star $\sqrt{}$star
External Knowledge Enhancement Representation Enhancement KnowMMR [25] IJCAI 2020 $\sqrt{}$ $\sqrt{}$ $\sqrt{}$
DSET [98] DOCS 2023 $\sqrt{}$ $\sqrt{}$star $\sqrt{}$
ECIFF [96] ICTAI 2023 $\sqrt{}$star $\sqrt{}$ $\sqrt{}$
ECIHS [100] INTERSPEECH 2023 $\sqrt{}$star $\sqrt{}$ $\sqrt{}$
KLPWE [102] Machine Learning 2024 $\sqrt{}$star $\sqrt{}$ $\sqrt{}$
DFP [99] COLING 2024 $\star$ $\star$ $\star$ $\sqrt{}$
Data Augmentation Knowdis [91] COLING 2020 $\sqrt{}$star $\sqrt{}$
CauSeRL [94] ACL-IJCNLP 2021 $\sqrt{}$star $\star$ ×
LearnDA [93] ACL-IJCNLP 2021 $\sqrt{}$star $\star$ ×
Reasoning Enhancement LSIN [103] ACL 2021 $\sqrt{}$ $\sqrt{}$star ×
Chen and Mao [105] ESWA 2024 $\star$ $\star$ $\star$ $\sqrt{}$
GCKAN [108] IS 2024 $\sqrt{}$star $\sqrt{}$ $\sqrt{}$
C3NET [107] KBS 2024 $\star$ $\sqrt{}$star $\sqrt{}$
DECI Event Graph Reasoning DCGIM [109] IS 2021 $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$
RichGCN [110] NAACL 2021 $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$
DocECI [114] APWeb-WAIM 2022 $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$
ERGO [112] COLING 2022 $\star$ $\sqrt{}$star - $\sqrt{}$
GESI [111] SIGIR 2022 $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$
CHEER [113] ACL 2023 $\star$ $\sqrt{}$star - $\sqrt{}$
PPAT [115] IJCAI 2023 $\star$ $\sqrt{}$star - $\sqrt{}$
GIMC [117] COLING 2024 $\sqrt{}$star $\star$ - $\sqrt{}$
iLIF [116] ACL 2024 $\star$ $\sqrt{}$star - $\star$
Prompt-based Fine-Tuning CPATT [118] KBS 2023 $\star$ $\sqrt{}$star - $\star$
DAPrompt [119] ACL 2023 $\star$ $\sqrt{}$star - $\sqrt{}$
HOTECI [120] COLING 2024 $\star$ $\sqrt{}$star - $\sqrt{}$
Other Strategies DPLSTM [121] ACL 2017 $\sqrt{}$ $\sqrt{}$ $\sqrt{}$
LR+ & LIP [122] NAACL 2019 × × $\sqrt{}$ $\sqrt{}$
KADE [123] AAAI 2023 $\star$ $\sqrt{}$star $\sqrt{}$
SENDIR [124] ACL 2023 $\star$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$
DiffusECI [125] AAAI 2024 $\star$ $\sqrt{}$star - $\star$
5 Experimental Evaluation
To evaluate the performance of current ECI models, we conducted comparative analyses of several state-of-the-art models using two commonly used datasets: Causal-TimeBank (CTB) and Event StoryLine v0.9 (ESL).
TABLE VI: Experimental Results in CTB. All the results have been multiplied by 100, the bold values represent the optimal results, while the underlined values indicate the suboptimal results.
Type Model Intra-sentence
P R F1
Feature Pattern-Based Matching RB 36.8 12.3 18.4
DD 67.3 22.6 33.9
VR-C 69.0 31.5 43.2
Deep Semantic Encoding Simple Encoding LSTM 18.3 43.5 25.8
BiLSTM 21.5 46.7 29.4
BERT 41.4 45.8 43.5
RoBERTa 39.9 60.9 48.2
Longformer 29.6 68.6 41.4
Enhanced Encoding SemSIn 52.3 65.8 58.3
ECLEP 50.6 63.4 56.3
CF-ECI 50.5 59.9 54.8
SemDI 59.3 77.8 67.0
Causal Knowledge Pre-training and Prompt-based Fine-tuning Causal Knowledge PLM Unified-QA 19.7 43.5 27.1
LLaMA2-7B 5.4 53.9 9.8
davinci-002 5.0 75.2 9.3
davinci-003 8.5 64.4 15.0
GPT-3.5-turbo 7.0 82.6 12.8
GPT-4.0 6.1 97.4 11.5
Prompt-based Fine-Tuning for PLM DPJL 63.6 66.7 64.6
GENECI 60.1 53.3 56.5
KEPT 48.2 60 53.5
GenSORL 60.1 53.3 56.3
ICCL 63.7 68.8 65.4
External Knowledge Enhancement Representation Enhancement KnowMMR 36.6 55.6 44.1
DSET 65.8 72.1 68.5
ECIFF 53.1 59.4 56.1
ECIHS 61.4 64.6 63.0
DFP 63.7 64.2 58.5
Data Augmentation CauSeRL 43.6 68.1 53.2
Knowdis 42.3 60.5 49.8
LearnDA 41.9 68.0 51.9
Reasoning Enhancement C3NET 60.2 72.8 65.9
Chen and Mao 83.7 55.5 65.2
GCKAN 52.2 60.7 56.1
LSIN 51.5 56.2 52.9
DECI Models Event Graph Reasoning CHEER 56.4 69.5 62.3
ERGO 62.1 61.3 61.7
iLIF 69.9 46.7 56.0
PPAT 67.9 64.6 66.2
RichGCN 39.7 39.7 46.7
Prompt-based Fine-Tuning CPATT 77.5 73.2 75.2
DAPrompt 66.3 67.1 65.9
HOTECI 71.1 65.9 68.4
Other Strategies DiffusECI 87.7 66.1 75.4
KADE 56.8 70.6 66.7
SENDIR 65.2 57.7 61.2
TABLE VII: Experimental Results in ESL. All the results have been multiplied by 100, the bold values represent the optimal results, while the underlined values indicate the suboptimal results.
Type Model Intra-sentence Inter-sentence Overall
P R F1 P R F1 P R F1
Feature Pattern -Based Matching OP 22.5 98.6 36.6 8.4 99.5 15.6 10.5 99.2 19.0
Deep Semantic Encoding Simple Encoding LSTM 34.0 41.5 37.4 13.5 30.3 18.7 17.6 33.9 23.2
BiLSTM 32.7 44.9 37.8 11.3 29.5 16.4 15.5 34.3 21.4
BERT 60.4 45.7 52.0 30.6 39.1 34.3 37.2 41.2 39.1
RoBERTa 62.7 45.4 52.7 32.7 38.3 35.3 39.7 40.6 40.1
Longformer 47.7 69.3 56.5 26.1 55.6 35.5 31.4 60.0 41.2
Enhanced Encoding SemSIn 64.2 65.7 64.9 - - - - - -
ECLEP 49.3 68.1 57.1 - - - - - -
CF-ECI 47.1 66.4 55.1 - - - - - -
SemDI 56.7 68.6 62.0 - - - - - -
Causal Knowledge Pre-training and Prompt-based Fine-tuning Causal Knowledge PLM Unified-QA 21.3 46.2 29.2 17.1 26.4 20.8 17.4 30.7 22.2
LLaMA2-7B 12.1 50.7 19.5 6.5 51.5 11.5 7.8 50.8 13.5
davinci-002 23.2 80.0 36.0 11.4 58.4 19.1 14.1 66.3 23.2
davinci-003 33.2 74.4 45.9 12.7 54.3 20.6 17.6 59.2 27.1
GPT-3.5-turbo 27.6 80.2 41.0 15.2 61.8 24.4 20.7 67.5 31.7
GPT-4.0 27.2 94.7 42.2 16.9 64.7 26.8 20.9 73.4 32.5
Prompt-based Fine-Tuning for PLM DPJL 65.3 70.8 67.9 - - - - - -
GENECI 59.5 57.1 58.8 - - - - - -
KEPT 50.0 68.8 57.9 - - - - - -
GenSORL 65.6 63.3 64.6 - - - - - -
ICCL 62.6 66.1 64.2 - - - - - -
External Knowledge Enhancement Representation Enhancement KnowMMR 41.9 62.5 50.1 - - - - - -
DSET 67.4 71.8 69.5 - - - - - -
ECIFF 51.9 62.8 56.8 - - - - - -
ECIHS - - 65.2 - - - - - -
DFP 55.9 69.8 62.1 - - - - - -
Data Augmentation CauSeRL 43.6 68.1 53.2 - - - - - -
Knowdis 42.3 60.5 49.8 - - - - - -
LearnDA 41.9 68.0 51.9 - - - - - -
Reasoning Enhancement C3NET 60.5 73.6 66.4 - - - - - -
Chen and Mao 67.4 69.5 68.3 - - - - - -
GCKAN 77.6 50.9 60.6 - - - - - -
LSIN 49.7 58.1 52.5 - - - - - -
DECI Models Event Graph Reasoning CHEER 56.9 69.6 62.6 45.2 52.1 48.4 49.7 53.3 51.4
DocECI 45.8 67.9 54.7 44.3 59.5 50.8 45.3 62.5 52.5
ERGO 57.5 72.0 63.9 51.6 43.3 47.1 48.6 53.4 50.9
GESI - - 50.3 - - 49.3 - - 49.4
iLIF 76.8 66.3 71.2 53.5 65.9 59.1 59.2 66.1 62.5
PPAT 62.1 68.8 65.3 54.0 50.2 52.0 56.8 56.0 56.4
RichGCN 49.2 63.0 55.2 39.2 45.7 42.2 42.6 51.3 46.6
Prompt-based Fine-Tuning CPATT 79.4 81.3 80.4 74.9 60.1 66.7 76.5 66.2 71.0
DAPrompt 64.5 73.6 68.5 59.9 59.3 59.0 61.4 63.7 62.1
HOTECI 66.1 72.3 69.1 81.4 40.6 55.1 63.1 51.2 56.5
Other Strategies LIP 38.8 52.4 44.6 35.1 48.2 40.6 36.2 49.5 41.9
LR+ 37.0 45.2 40.7 25.2 48.1 33.1 27.9 47.2 35.1
DiffusECI 65.3 78.3 71.4 61.9 59.9 60.9 63 64.1 63.5
KADE 61.5 73.2 66.8 51.2 74.2 60.5 51.9 70.6 59.8
SENDIR 65.8 66.7 66.2 33 90 48.3 37.8 82.8 51.9
(a) CTB (Intra-Sentence)
(b) ESL (Intra-Sentence)
(c) ESL (Inter-Sentence)
(d) ESL (Overall)
Figure 12: Comparison of P, R, and F1 across different model types on CTB and ESL datasets. The error bars represent the range of variability for each metric across different models. Each error bar shows the maximum and minimum values observed for a given metric within a specific model type.
5.1 Experimental Setup
5.1.1 Dataset Processing
The details of the two datasets used in this experiment, Causal-TimeBank (CTB) and Event StoryLine v0.9 (ESL), are provided in Section 2.2. For consistency with prior studies, we adopted the common data splits: in the ESL dataset, the last two topics were designated as development data, while the remaining 20 topics were used for 5-fold cross-validation. For the CTB dataset, we conducted 10-fold cross-validation.
5.1.2 Evaluation Protocol
We evaluated the models using Precision (P), Recall (R), and F1-score (F1) as metrics, distinguishing between intra-sentence causality, inter-sentence causality, and overall performance. Given that most existing methods evaluate only the presence of causalities without considering directional accuracy, we focused our comparison similarly on causality presence alone. For the CTB dataset, where only 18 inter-sentence causal pairs exist, we limited evaluation to intra-sentence causality. For the ESL dataset, we assessed performance on intra-sentence causality, inter-sentence causality, and overall identification.
5.1.3 Baseline Models
We included several state-of-the-art models from existing literature, along with standard baselines like LSTM and BERT (used as encoding and classification frameworks). Additionally, we incorporated multiple LLMs as baselines, including GPT-3 (text-davinci), GPT-3.5-turbo, GPT-$4$https://openai.com/chatgpt/, and LLaMA2-7B  [^5] http://huggingface.co/meta-llama/Llmma-2-7b. For the LLMs, we used a standardized prompt:
[D] [SEP] Is there a causal relationship between [E1] and [E2]? Think step-by-step to get the answer.
Here, if [E1] and [E2] are within the same sentence, D represents that sentence; otherwise, D represents the entire document. It should be noted that many DECI models also support SECI, so models specifically designed for DECI are included in SECI performance comparisons where applicable. The open-source code for some models is released on $\mathtt{Gitee}$ [^6] https://gitee.com/KGAWS/eci-related-works-and-datasets.
5.1.4 Implementation Details
To ensure fair comparisons, if baseline models had reported results on CTB and ESL with consistent data splits, we directly cited those results. For baselines where the original paper used different datasets or splits but provided open-source code, we re-evaluated the model by reproducing it according to our specified data splits. Reproduction strictly followed hyperparameters reported in the original paper or code to maintain result comparability. Models without open-source code or using entirely different datasets were excluded. Additionally, for models with multiple variants, we reported the variant with the highest F1-score.
5.2 Results and Analysis
Tables VI and VII present the results of baseline methods on the CTB and ESL datasets, respectively.
On the CTB dataset, DiffusECI achieved the best overall performance (F1=75.4%), followed by CPATT with a close second-best overall performance (F1=75.2%). DiffusECI also attained the highest precision with the fewest false positives (P=87.7%), while the model by Chen and Mao ranked second in precision (P=83.7%). In terms of recall, GPT-4.0 achieved the highest score (R=97.4%), with GPT-3.5-turbo as the second-best (R=82.6%).
On the ESL dataset, for intra-sentence causality, CPATT achieved the highest overall performance (F1=80.4%), with DiffusECI as the second-best (F1=71.4%). CPATT also recorded the highest precision (P=79.4%), while GCKAN followed with a close second (P=77.6%). For recall, OP had the highest score (R=98.6%), followed by GPT-4.0 (R=94.7%). For inter-sentence causality on the ESL dataset, CPATT again achieved the best overall performance (F1=66.7%), with DiffusECI as the second-best (F1=60.9%). HOTECI demonstrated the highest precision (P=81.4%), followed by CPATT (P=74.9%). In terms of recall, OP led with the highest score (R=99.5%), with SENDIR as the second-best (R=90%). For overall causality, CPATT achieved the best overall performance (F1=71.0%), with DiffusECI following as the second-best (F1=63.5%). CPATT also achieved the highest precision (P=76.5%), with HOTECI as the runner-up (P=63.1%). In terms of recall, OP reached the highest score (R=99.2%), followed by SENDIR (R=82.8%). Overall, CPATT and DiffusECI demonstrated the strongest performances among all baselines across various settings.
To analyze the performance of different types of methods, we calculated the average performance for each method category. Figure 12 displays the comparative results across these method types.
5.2.1 Baseline Performance on SECI Task
Feature pattern matching-based methods struggle with balancing precision and recall, resulting in limited overall performance. Shallow feature-based methods, such as DD and VR-C, show high precision but low recall, while simple statistical methods like PPMI underperform in both.
Deep semantic encoding-based methods improve recall but produces many false positives, likely due to a limited understanding of "causality" versus "correlation". Enhanced encoding-based methods show notable improvements, benefiting from richer features and better event context representation.
CKPLM-based methods achieve high recall but with frequent false positives, a result of inherent "causal hallucination" in PLMs/LLMs. The performance of these methods also varies significantly with model scale. Prompt-based fine-tuning methods, by contrast, offer balanced performance, achieving solid precision and recall across SECI tasks in both datasets.
External knowledge improves results over basic semantic encoding, showing that it effectively boosts ECI performance. Among these approaches, reasoning- and representation-enhanced methods show more balanced performance, especially excelling in precision compared to data augmentation-based methods. Although data augmentation-based methods increase recall over shallow and simple encoding methods, they also introduce more false positives, indicating that the quality of additional training data is crucial for enhancing overall performance.
Event graph reasoning, designed for DECI tasks, also performs well on SECI tasks with balanced precision and recall, only slightly behind PLM prompt-tuning and reasoning-enhanced methods. Some models, like iLIF and RichGCN, still face challenges with balancing precision and recall due to design nuances.
Prompts-based fine-tuning methods achieve a favorable balance between precision and recall, resulting in optimal overall performance. It significantly outperforms other baselines in terms of precision, indicating its capability in suppressing false positives. This advantage is primarily attributed to the effective constraints provided by the prompts.
5.2.2 Baseline Performance on DECI Task
Feature pattern recognition methods, such as OP, LIP, and LR+, deliver relatively high recall but with many false positives. Simple encoding-based methods show low precision and recall due to difficulty capturing context and causal links, especially between distant events.
CKPLM-based methods perform similarly on DECI and SECI tasks, with high recall but lower precision. However, recall decreases significantly in DECI tasks, suggesting that even advanced PLMs struggle to accurately identify inter-sentence causal relations in unsupervised settings.
Prompt-based fine-tuning methods again achieve the best-balanced overall performance, particularly excelling in precision by reducing false positives, indicating that high-quality and effective prompts substantially improve PLM understanding of causal patterns.
Event graph reasoning-based methods also perform well, showing higher recall but slightly lower precision. This drop is likely due to directly modeling events as graph nodes, potentially missing textual details and increasing false positives. However, the method’s higher recall results from effectively reducing noise, improving representation quality, and capturing causal links across distant events.
6 Future Directions
In the previous sections, we have reviewed the recent advances in ECI. Although more and more advanced models have been introduced to enhance event representation and causal reasoning, here are still many challenges and open problems that need to be addressed. In this section, we discuss the future directions of this research area.
6.1 Causal Direction Identification
Current ECI research primarily focuses on determining the existence of causalities, often overlooking the direction of causality. However, causal direction is crucial for information extraction and knowledge reasoning, especially in constructing a comprehensive event timeline to inform analysis and decision-making [116]. Identifying causal direction requires a deeper semantic understanding and representation learning. Future research could explore directed graph-based inference methods, potentially integrating PLMs, GNNs, and causal inference techniques to enhance directional reasoning [126].
6.2 Event Causal Chain Reasoning
Identified causal relations can be linked to form causal chains, supporting decision-making and providing insights into event development. Understanding these chains is essential for tracing key events and analyzing the sequence of occurrences. However, simply concatenating causal event pairs may be inadequate due to issues like context drift and threshold effects [127]. Future research should investigate causal chain reasoning, leveraging traditional causal inference methods such as counterfactual data augmentation [128] and Structural Causal Model (SCM) [129] to build reliable causal chains.
6.3 Uncertain ECI
In real-world scenarios, causality often involves degrees of uncertainty, requiring a "confidence level" or "conditional" reasoning approach. This introduces challenges in implementing conditional and fuzzy reasoning [130] within ECI frameworks. Additionally, existing labeled datasets lack confidence information, typically providing only binary or ternary causal labels. Future efforts should focus on developing datasets that include confidence annotations, despite the complex labeling process, as these are essential for real-world applications.
6.4 Event Causality Extraction
ECI typically assumes given event mentions, but identifying relevant events in text first requires event extraction (EE) [1]. Most studies use EE algorithms to extract events and trigger words before applying ECI, which can introduce error propagation. A more effective approach may involve simultaneous event extraction and causal identification, termed Event Causality Extraction (ECE) or Event Causality Generation (ECG) [131]. ECE is more complex than ECI but offers significant potential, as event classification and argument extraction in EE help clarify semantic and logical connections between events, thus supporting ECI.
6.5 Multi-Lingual ECI
Multi-lingual ECI has recently gained attention. While current models perform well in monolingual settings, they often lack generalization across languages, highlighting the need for robust multilingual models. Techniques like language-independent feature extraction, transfer learning, machine translation, meta-learning, and multilingual PLMs hold promise for multi-lingual ECI (MECI) tasks [132]. The multilingual dataset published by Lai et al. [46] provides a strong foundation for future MECI research.
6.6 Multimodal ECI
In most ECI tasks, events and contextual information are derived solely from text. However, with advancements in multimodal technologies, events might originate from images or videos, with additional context provided by these sources. Leveraging multimodal data could significantly enhance causal reasoning [133] but requires entity disambiguation across modalities and alignment of semantic representations using multimodal pre-trained models. Developing multimodal ECI datasets is also crucial, potentially achieved by refining existing multimodal relation extraction datasets [134, 135].
6.7 Explainable ECI
Explainability is vital in causal inference research. Beyond detecting causalities, understanding the mechanisms and reasoning paths behind them is essential. Current state-of-the-art methods rely heavily on "black-box" deep neural networks, which lack transparency and reduce user trust [136]. Future work should focus on developing methods that provide interpretable causal explanations, clarifying the reasoning steps and evidence supporting causal conclusions. Enhanced explainability will improve model reliability, transparency, and applicability to real-world tasks [137].
6.8 Few-Shot and Zero-Shot ECI
While some methods [27, 86] exhibit few-shot and zero-shot ECI capabilities, they tend to generate high false positives, limiting practical applications. Moreover, performance drops significantly in DECI tasks under few-shot or zero-shot conditions. Future research could address these issues by using LLM-based data augmentation on existing labeled datasets [138] to generate diverse causal syntactic and grammatical patterns, or by converting text to semantic graph formats to constrain LLM outputs [139], thereby mitigating causal hallucination problems.